State Cramer’s rule. Show graphically how a system of two equations in 

two unknowns has:

(i) no solution, (ii) a unique, (iii) infinitely many solutions.


1
Expert's answer
2021-03-30T16:40:17-0400

Cramer’s rule:

The solution of the simultaneous equations

"\begin{cases}
a_1x+b_1y=d_1
\\
a_2x+b_2y=d_2

\end{cases}"

is given by the formulae

"x=\frac{\begin{vmatrix}
 d_1 & b_1 \\
 d_2 & b_2
\end{vmatrix}

}{\begin{vmatrix}
 a_1 & b_1 \\
 a_2 & b_2
\end{vmatrix}}" "y=\frac{\begin{vmatrix}
 a_1 & d_1 \\
 a_2 & d_2
\end{vmatrix}

}{\begin{vmatrix}
 a_1 & b_1 \\
 a_2 & b_2
\end{vmatrix}}"


(i) "\begin{cases}
x-y=0
\\
x-y=1

\end{cases}"

This system has no solution.





ii)"\begin{cases}
x-y=0
\\
x+y=0

\end{cases}" 

This system has a unique solution.



(iii) "\begin{cases}
x-y=1
\\
2x-2y=2
\end{cases}" 

This system has infinitely many solutions.